Please help, no text book, online class. Thanks in advance, you rock!!!
1. What is the variable used to predict the value of another called?
A) Independent
B) Dependent
C) Correlation
D) Determination


2. Based on the regression equation, we can
A) predict the value of the dependent variable given a value of the independent variable.
B) predict the value of the independent variable given a value of the dependent variable.
C) measure the association between two variables.


3. What is the range of values for a coefficient of correlation?
A) 0 to +1.0
B) –3 to +3 inclusive
C) –1.0 to +1.0 inclusive
D) Unlimited range


4. What is the general form of the regression equation?
A) Y' = ab
B) Y' = a + bX
C) Y' = a – bX
D) Y' = abX



5. Which value of r indicates a stronger correlation than 0.40?
A) –0.30
B) –0.50
C) +0.38
D) 0


6. Suppose the least squares regression equation is Y' = 1202 + 1133X.
When X = 3, what does Y' equal?
A) 5,734
B) 8,000
C) 4,601
D) 4,050


7. In the least squares equation, Y' = 10 + 20X , the value of 20 indicates
A) the Y intercept.
B) for each unit increase in X, Y increases by 20.
C) for each unit increase in Y, X increases by 20.


8. In the equation Y' = a + bX, what is Y'?
A) Slope of the line
B) Y intercept
C) Predicted value of Y, given a specific X value
D) Value of Y when X=0



9. Based on the regression equation, we can
A) predict the value of the dependent variable given a value of the independent variable.
B) predict the value of the independent variable given a value of the dependent variable.
C) measure the association between two variables.


10. If the regression equation is Y' = 2 – 0.4X, what is the value of Y' when X = –3?
A) 0.8
B) 3.2
C) –10.0
D) 14.0


11. In the regression equation, y’ = a + bx what does the letter "b" represent?
A) Y intercept
B) Slope of the line
C) Any value of the independent variable that is selected
D) Value of Y when X=0


12. A sales manager for an advertising agency believes there is a relationship between the number of contacts and the amount of the sales. What is the dependent variable?
A) Salesperson
B) Number of contacts
C) Amount of sales
D) All the above

Answer by stanbon(57387)

1. What is the variable used to predict the value of another called?
A) Independent
B) Dependent
C) Correlation
D) Determination
Ans: Independent
-------------------------------------------
2. Based on the regression equation, we can
A) predict the value of the dependent variable given a value of the independent variable.
B) predict the value of the independent variable given a value of the dependent variable.
C) measure the association between two variables.
Ans: A
---------------------------------------------
3. What is the range of values for a coefficient of correlation?
A) 0 to +1.0
B) –3 to +3 inclusive
C) –1.0 to +1.0 inclusive
D) Unlimited range
Ans: C
-------------------------
4. What is the general form of the regression equation?
A) Y' = ab
B) Y' = a + bX
C) Y' = a – bX
D) Y' = abX
Ans: B
-----------------

5. Which value of r indicates a stronger correlation than 0.40?
A) –0.30
B) –0.50
C) +0.38
D) 0
Ans B
----------------------------------
6. Suppose the least squares regression equation is Y' = 1202 + 1133X.
When X = 3, what does Y' equal?
A) 5,734
B) 8,000
C) 4,601
D) 4,050
Ans : C
----------------
7. In the least squares equation, Y' = 10 + 20X , the value of 20 indicates
A) the Y intercept.
B) for each unit increase in X, Y increases by 20.
C) for each unit increase in Y, X increases by 20.
Ans B
------------
8. In the equation Y' = a + bX, what is Y'?
A) Slope of the line
B) Y intercept
C) Predicted value of Y, given a specific X value
D) Value of Y when X=0
Ans C
---------------------------------------------
9. Based on the regression equation, we can
A) predict the value of the dependent variable given a value of the independent variable.
B) predict the value of the independent variable given a value of the dependent variable.
C) measure the association between two variables.
Ans: A
----------------------
10. If the regression equation is Y' = 2 – 0.4X, what is the value of Y' when X = –3?
A) 0.8
B) 3.2
C) –10.0
D) 14.0
Ans: C
--------------

11. In the regression equation, y’ = a + bx what does the letter "b" represent?
A) Y intercept
B) Slope of the line
C) Any value of the independent variable that is selected
D) Value of Y when X=0
Ans: B
---------------

12. A sales manager for an advertising agency believes there is a relationship between the number of contacts and the amount of the sales. What is the dependent variable?
A) Salesperson
B) Number of contacts
C) Amount of sales
D) All the above
Ans: C
---------------
cheers,
Stan H.